the sides of a triangle are in the proportion 2:3:7. the longest side of the triangle is 14 cm. What is the perimeter of the triangle?
Answer by stanbon(75887)

the sides of a triangle are in the proportion 2:3:7. the longest side of the triangle is 14 cm. What is the perimeter of the triangle?
------
2:3:7 is the same as 2x:3x:7x
-----
If 7x = 14, x = 2
-----
Perimeter = 2x+3x+7x = 12x = 12*2 = 24
---------------------
Cheers,
Stan H.
===============